The speed of a car is reduced from 30m/s in a distance of 100m on applying the brakes assuming the retardation to be uniform, find the time taken (b). how much further will the car travel before coming to rest
1
Expert's answer
2018-10-31T13:57:09-0400

Distance traveled by car

s=(v_i+v_f)/2 t

So, the time of motion to the rest

t=2s/(v_i+v_f )=(2×100 m)/(30 m/s+0 m/s)=6.67 s